X) The answers: a) 0.315 V c) 0.273 V d) 1.80 M10. For the voltaic cell: Cd(s) Cd2 (0.100 M, 25 mL) || Pb2 (2.000 M, 25 mL) | Pb(s) a. What is Eel initially? (E 0.277 v) b.
Answer #1

10)

a)

Cd(s) + Pb2+(aq) ------->   Cd2+(aq)   +   Pb(s)

E0cell = E0(Pb2+/Pb) - E0(Cd2+/Cd)

E0cell =    0.277 V

Ecell = E0cell + ( 0.059 / n ) log ( Pb2+ / Cd2+)

Ecell = 0.277 + ( 0.059 / 2 ) * log ( 2 / 0.1)

Ecell = 0.277 + ( 0.059 / 2 ) * log ( 20 )

Ecell = 0.277 +   0.038

Ecell = 0.315 V

so Ecell = 0.315 V

b)

during the operation the Cd will oxidize to Cd2+ and Pb2+ will reduce to Pb

so the concentration of Pb2+ will decrease and Cd2+ will increase

log ( Pb2+ / Cd2+)  will decrease and hence the Ecell will decrease during the operation of cell

c)

when [ Pb2+] = 0.9 M

so change in [Pb2+] = 2 - 0.9 = 1.1 M

since both oxidation of Cd to Cd2+ and reduction Pb2+ to Pb will happen

so the concentration of Cd2+ will change corresponding to concentration change in Pb2+

so the concentration of Cd2+ increase by 1.1 M

[ Cd2+ ] = 0.1 + 1.1 = 1.2 M

Ecell = E0cell + ( 0.059 / n ) log ( Pb2+ / Cd2+)

Ecell = 0.277 + ( 0.059 /  2 ) log ( 0.9 / 1.2 )

Ecell = 0.277 + ( 0.059 /  2 ) log ( 0.75)

Ecell = 0.277 - 0.004

Ecell = 0.273 V

Ecell = 0.273 V

d)

Ecell = E0cell + ( 0.059 / n ) log ( Pb2+ / Cd2+)

0.254   = 0.277 + ( 0.059 /  2 ) log ( Pb2+ / Cd2+)

0.254 - 0.277 = 0.0295 * log ( Pb2+ / Cd2+)

-0.023 / 0.0295 = log ( Pb2+ / Cd2+)

- 0.7797   = log ( Pb2+ / Cd2+)

Pb2+ / Cd2+ = 10 ^ ( - 0.7797 )

Pb2+ / Cd2+   = 0.166

Pb2+ = 0.166 * Cd2+

at any time

[ Pb2+] + [Cd2+] = 2.1

0.166 * Cd2+    +   Cd2+   = 2.1

1.166 Cd2+ = 2.1

[ Cd2+] = 2.1 / 1.166

[Cd2+] = 1.80 M

so the [ Cd2+ ] = 1.80 M     At Ecell = 0.254 V